One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by mszlmb(115)

One number is 12 more than another.  The sum of the smaller number and twice the 
larger number is 39. Find the larger number. Let us assign x to the smaller number and y to the larger number. From the first sentence we derive the following equation: X=Y+12 From the second, we derive the following: Y+2X=39 We know X is equal to Y+12, so we'll replace the X in the second equation with (Y+12). The second equation now yields: Y+2(Y+12)=39. Let us simplify by distribution: Y+2Y+24=39 Simplify and minus 24 on both sides: 3Y=15 Divide both sides by 5: Y=5 We found Y! Since we know Y is equal to 5, we can replace it into the equation X=Y+12.
This yields: X=5+12, or 17. X=17 Y=5